# CancerGUIDE: Cancer Guideline Understanding via Internal Disagreement Estimation

### v1

@@ Abstract
The National Comprehensive Cancer Network (NCCN) provides evidence-based guidelines for cancer treatment. Translating complex patient presentations into guideline-compliant treatment recommendations is time-intensive, requires specialized expertise, and is prone to error. Advances in large language model (LLM) capabilities promise to reduce the time required to generate treatment recommendations and improve accuracy. We present an LLM agent-based approach to automatically generate guideline-concordant treatment trajectories for patients with non-small cell lung cancer (NSCLC). Our contributions are threefold. First, we construct a novel longitudinal dataset of 121 cases of NSCLC patients that includes clinical encounters, diagnostic results, and medical histories, each expertly annotated with the corresponding NCCN guideline trajectories by board-certified oncologists. Second, we demonstrate that existing LLMs possess domain-specific knowledge that enables high-quality proxy benchmark generation for both model development and evaluation, achieving strong correlation (Spearman coefficient $r=0.88$, RMSE = $0.08$) with expert-annotated benchmarks. Third, we develop a hybrid approach combining expensive human annotations with model consistency information to create both the agent framework that predicts the relevant guidelines for a patient, as well as a meta-classifier that verifies prediction accuracy with calibrated confidence scores for treatment recommendations (AUROC=0.800), a critical capability for communicating the accuracy of outputs, custom-tailoring tradeoffs in performance, and supporting regulatory compliance. This work establishes a framework for clinically viable LLM-based guideline adherence systems that balance accuracy, interpretability, and regulatory requirements while reducing annotation costs, providing a scalable pathway toward automated clinical decision support.
@@
We obtain clinical annotations from a third-party contractor. Patient data is de-identified and sourced from multiple institutions within the USA, but it is not licensed for public release. Code is made available here: CancerGUIDE repository.
@@ 1 Introduction
Cancer treatment decisions require oncologists to synthesize complex patient histories with evolving clinical guidelines to recommend the appropriate next treatments. The National Comprehensive Cancer Network (NCCN) guidelines provide evidence- and consensus-based recommendations for cancer diagnosis, treatment, and management (National Comprehensive Cancer Network, 2025), with adherence promoting higher quality and consistent care between providers (Pluchino and D’Amico, 2020; Winn and McClure, 2003; Benson and Brown, 2008). However, guideline navigation presents significant challenges: the guidelines are extensive, frequently updated as new research emerges, and require time-intensive review of complex patient documentation (Benson and Brown, 2008). These factors contribute to variability in guideline adherence and treatment recommendations, particularly in resource-constrained settings where specialist expertise is limited (Koh et al., 2020; Kumar et al., 2023).
Large language models (LLMs) offer promising potential to address these challenges by automatically processing clinical notes and recommending guideline-concordant treatment plans (Fast et al., 2024; Wang et al., 2024; Hamed et al., 2023; Schulte, 2023). However, deploying LLMs for clinical decision support requires rigorous evaluation to ensure accuracy and safety. The complexity of clinical reasoning, combined with the high stakes of treatment decisions, demands robust validation methods that can assess model performance at scale while maintaining clinical safety standards. In addition, current FDA guidelines for evaluation of AI systems recommend ROC curve measurement as a part of comprehensive clinical performance assessment endpoints (U.S. Food and Drug Administration, 2022), which is challenging to produce from generative outputs that are not associated with semantically aligned confidence scores.
Despite growing interest in clinical LLMs, rigorous evaluation remains a fundamental bottleneck due to the scarcity of expert-annotated datasets. High-quality ground truth labels for complex clinical reasoning tasks require substantial investment in specialist time and expertise, limiting the scale at which models can be validated (Zajac et al., 2023; Liang et al., 2017). Common evaluation approaches face significant limitations: synthetic data generation often fails to capture clinical complexity and is vulnerable to distributional shift (Santangelo et al., 2025; Yan et al., 2022; Chen et al., 2024; Cui et al., 2025), while using actual patient treatments as ground truth is problematic, since real-world decisions frequently incorporate factors exogenous to guideline recommendations, such as patient preferences, drug availability, institutional protocols, and physician experience (Arts et al., 2016; Quaglini et al., 2004).
@@
Figure 1: CancerGUIDE framework: (1) Clinicians derive patient pathways ($label_{doc}$) from NCCN guidelines and real notes, which are then compared with LLM-derived pathways ($label_{LLM}$) to calculate $Score_{r}$. (2) Proxy benchmarks use NCCN guidelines and both synthetic and real clinical notes to evaluate LLMs by comparing their outputs ($label_{LLM}$) against weak labels ($label_{proxy}$) to calculate $Score_{p}$ and measure similarity of performance on expert annotations. (3) A consistency framework uses model consistency features (Feature Sets) and $Score_{p}$ in a meta-classifier to decide whether to accept or reject treatment predictions.
@@
Our evaluation highlights the role of synthetic data and consistency-based data in proxy benchmarking, with these approaches achieving high Spearman correlation coefficients (r=0.88) and low RMSE values (RMSE=0.08). We demonstrate that model consistency can also serve as a reliable accuracy predictor, enabling our meta-learning framework to achieve an average of 0.800 AUROC in classifying individual prediction accuracy across all models, while unsupervised clustering using consistency signals alone achieves 0.666 F1 at separating correct from incorrect predictions.
@@
    Consistency framework for reliable treatment prediction: A hybrid agent and meta-classifier framework that generates relevant guideline paths and produces confidence values correlated with accuracy, enabling ROC curve calculation to clearly convey performance to clinicians and ensure compliance with regulatory standards (U.S. Food and Drug Administration, 2022).
@@ 2.1 Guideline Adherent Treatment Recommendations
Guideline-adherent treatments have consistently been associated with improved clinical outcomes, such as overall survival (Dholakia et al., 2020; Lindqvist et al., 2022; Wöckel et al., 2010). Clinical decision-support systems have the ability to translate clinical guidelines into point-of-care recommendations. Several studies have demonstrated the use of AI-driven systems in predicting guideline-adherent treatment recommendations in oncology (Nafees et al., 2023; Suwanvecho et al., 2021). Emerging work has also explored the utility of LLMs; for example, preliminary investigations suggest that ChatGPT can effectively summarize guideline content (Hamed et al., 2023) and exhibits partial concordance in identifying guideline-adherent treatments (Schulte, 2023). Furthermore, recent benchmarking efforts have systematically evaluated the alignment of LLMs with established medical guidelines, underscoring both their potential utility and the current limitations imposed by the relatively small number of verified clinical cases available (Fast et al., 2024).
@@ 2.2 Synthetic Generation of Clinical Data
Synthetic data generation is a promising strategy to address the challenges of privacy and data scarcity in healthcare. Past works have focused on generating high-fidelity electronic health records (EHRs) using generative models, such as generative adversarial networks, variational autoencoders and autoregressive models (Choi et al., 2017; Lu et al., 2022; Li et al., 2023; Yoon et al., 2023; Jarrett et al., 2021). More recent frameworks have used LLMs to generate synthetic data and have shown improvements in privacy guarantees and scalability (Altalla’ et al., 2025; Barr et al., 2025; Taloni et al., 2025; Litake et al., 2024). In parallel, standardized evaluation metrics and scorecards have been proposed to assess the fidelity, privacy, and clinical utility of synthetic health data, providing a foundation for more rigorous benchmarking and deployment (Santangelo et al., 2025; Yan et al., 2022; Chen et al., 2024; Bedi et al., 2025). In addition, synthetic cohorts have been developed to model guideline-adherent treatment pathways, such as synthetic stroke registries for adherence benchmarking (Wang et al., 2025) and synthetic EHR modules that embed clinical practice guidelines and protocols (Walonoski et al., 2017). These efforts demonstrate the potential of synthetic data for adherence-focused tasks. However, significant challenges remain in generating high-fidelity, guideline-adherent cases while maintaining performance on downstream evaluation tasks.
@@ 2.3 Weak Supervision and Consistency
Weak supervision and consistency-driven learning have emerged as key strategies for training models when labeled data is limited or noisy. Weak supervision techniques, such as programmatic labeling and distant supervision, integrate heterogeneous, partially labeled sources to produce probabilistic labels for downstream predictive tasks (Mintz et al., 2009; Ratner et al., 2016, 2017; Hsu and Roberts, 2025; Saad-Falcon et al., 2025). Consistency-based methods encourage models to produce stable predictions under input perturbations, data augmentations, or repeated evaluations, helping to regularize training and improve generalization (Laine and Aila, 2016; Miyato et al., 2019; Xie et al., 2020). Pseudo-labeling, a related strategy, generates labels for unlabeled or partially labeled data by treating confident model predictions as additional supervised training data, enhancing model performance by expanding the training set with its own high-confidence predictions (Lee, 2013; Zou et al., 2021; Xie et al., 2020). Within healthcare, pseudo-labeling has shown to improve the reliability of EHR phenotyping and imaging tasks under label scarcity (Dong et al., 2023; Zhang et al., 2022; Nogues et al., 2022). Together, these techniques provide a framework for leveraging limited labeled data while maintaining prediction stability and adherence to domain-specific constraints, which is particularly relevant for guideline-adherent treatment modeling.
@@ 2.4 Non-Verifiable Task Evaluation
Evaluating LLMs on open-ended or non-verifiable tasks, such as treatment recommendation prediction, is challenging due to the scarcity of ground truth labels. Traditional metrics like accuracy or BLEU are insufficient in zero-label settings, prompting the use of human-aligned and/or proxy evaluation frameworks (Lyu et al., 2024; Yamauchi et al., 2025; Tan et al., 2024). Further, studies have shown that consistency-derived benchmarks provide improved predictive fidelity relative to ground-truth outcomes (Lee et al., 2024; Li et al., 2025). Proxy frameworks and consistency-derived benchmarks offer partial solutions, but their alignment with expert judgment remains imperfect, motivating the continued development of hybrid evaluation strategies.
@@ 3.1 Guideline-Compliance Task Formalization
To ground this formalization in clinical practice, we leveraged o3’s vision capabilities (OpenAI, 2025c) to extract the NCCN NSCLC guideline decision tree (National Comprehensive Cancer Network, 2025) and curated an expert-annotated dataset. Oncologists traced patient notes through the decision tree, recording node sequences and assessing guideline adherence for each case. Further details on annotation procedures, participant recruitment, and quality control measures are provided in Appendix A.
@@
where $\mathcal{D}$ denotes the distribution over patients and $\mathcal{L}$ is standard loss. This formulation enables evaluation of $f$ in settings with limited or no access to human-labeled $(x,y)$ pairs, by leveraging mode agreement, self-consistency, and benchmark proxy-derived features as signals for meta-classification.
@@ 3.2.1 Synthetic Supervision
Two complementary strategies are used: Structured Generation fills empty structured fields conditioned on the generated path and full clinical guidelines, performs consistency checks by reconstructing implied paths (discarding mismatched cases), then generates unstructured notes from structured data, target path, guidelines, and real clinical note examples. Unstructured Generation bypasses structured fields, generating synthetic notes directly from target paths, guidelines, and clinical note examples.
@@ 3.2.2 Consistency-Based Pseudo-Labeling
For real clinical notes, pseudo-labels are assigned using model self-consistency:
  1. 1.
    Sample $k$ independent predictions $f(x)=\{\hat{y}_{1},\dots,\hat{y}_{k}\}$ from the same model.
  2. 2.
    Measure agreement along two axes: path overlap (structural alignment) and treatment match (final treatment concordance).
  3. 3.
    Retain notes with agreement above a threshold $\delta=0.9$; assign inconsistent cases score of 0.
@@
We further assign pseudo-labels by comparing predictions across models $M$. For a note $x$, if two or more models converge on the same pseudo-label after $k$ independent samples, we define this as the aggregated label. Inconsistent cases are simply excluded rather than penalized.
@@ 3.3 Final Treatment Accuracy Prediction
We use consistency signals to predict accuracy and identify errors without human labels. We train a meta-classifier using features derived from self-consistency metrics ($k$-rollout path overlap and treatment match), cross-model consistency (fraction of models with the same generated path), and proxy benchmark scores (synthetic and consistency-derived). This framework achieves an average of 0.800 AUROC in predicting whether the output of a model is correct, providing practitioners with a means of assessing confidence in model recommendations.
@@
To demonstrate label-free evaluation capabilities, we apply unsupervised methods to both accuracy classification and error identification. Clustering self- and cross-model consistency features naturally separates high-confidence correct from low-confidence incorrect predictions, achieving 0.666 F1 score without any human annotations. For error identification, we tabulate inconsistencies across $k$ rollouts for each model $m$ on patient $p$ and find that 40.42% of all model errors on human-labeled data can be detected without any human labels (detailed analysis in Appendix D). This demonstrates that consistency signals alone provide sufficient structure to both differentiate between accurate and inaccurate predictions and identify specific failure modes, enabling model developers to detect and address potential issues prior to deployment for iterative refinement of clinical decision support systems.

### v2

@@ Abstract
The National Comprehensive Cancer Network (NCCN) provides evidence-based guidelines for cancer treatment. Translating complex patient presentations into guideline-compliant treatment recommendations is time-intensive, requires specialized expertise, and is prone to error. Advances in large language model (LLM) capabilities promise to reduce the time required to generate treatment recommendations and improve accuracy. We present an LLM agent-based approach to automatically generate guideline-concordant treatment trajectories for patients with non-small cell lung cancer (NSCLC). Our contributions are threefold. First, we construct a novel longitudinal dataset of 121 cases of NSCLC patients that includes clinical encounters, diagnostic results, and medical histories, each expertly annotated with the corresponding NCCN guideline trajectories by board-certified oncologists. Second, we demonstrate that existing LLMs possess domain-specific knowledge that enables high-quality proxy benchmark generation for both model development and evaluation, achieving strong correlation (Spearman coefficient $r=0.88$, RMSE = $0.08$) with expert-annotated benchmarks. Third, we develop a hybrid approach combining expensive human annotations with model consistency information to create both the agent framework that predicts the relevant guidelines for a patient, as well as a meta-classifier that verifies prediction accuracy with calibrated confidence scores for treatment recommendations (AUROC=0.804). Calibrated confidence scoring is a critical capability for communicating the accuracy of outputs, custom-tailoring tradeoffs in performance, and supporting regulatory compliance. This work establishes a framework for clinically viable LLM-based guideline adherence systems that balance accuracy, interpretability, and regulatory requirements while reducing annotation costs, providing a scalable pathway toward automated clinical decision support.
@@
We obtain clinical annotations from a third-party contractor. Patient data is de-identified and sourced from multiple institutions within the USA, but it is not licensed for public release. All analyses were conducted using Azure-hosted LLMs within our institution’s secure cloud environment; no data were transmitted to external third-party services. Code and synthetic patient data are made available here: CancerGUIDE repository.
@@ 1 Introduction
Cancer treatment decisions require oncologists to synthesize complex patient histories with evolving clinical guidelines to recommend the appropriate next treatments. The National Comprehensive Cancer Network (NCCN) guidelines provide evidence- and consensus-based recommendations for cancer diagnosis, treatment, and management (NCCN), with adherence promoting higher quality and consistent care between providers (Pluchino2020; Winn2003; Benson2008). However, guideline navigation presents significant challenges: the guidelines are extensive, frequently updated as new research emerges, and require time-intensive review of complex patient documentation (Benson2008). These factors contribute to variability in guideline adherence and treatment recommendations, particularly in resource-constrained settings where specialist expertise is limited (Koh2020; Kumar2023).
Large language models (LLMs) offer promising potential to address these challenges by automatically processing clinical notes and recommending guideline-concordant treatment plans (Fast2024; Wang2024; Hamed2023; Schulte2023). However, deploying LLMs for clinical decision support requires rigorous evaluation to ensure accuracy and safety. The complexity of clinical reasoning, combined with the high stakes of treatment decisions, demands robust validation methods that can assess model performance at scale while maintaining clinical safety standards. In addition, current FDA guidelines for evaluation of AI systems recommend ROC curve measurement as a part of comprehensive clinical performance assessment endpoints (fda), which is challenging to produce from generative outputs that are not associated with semantically aligned confidence scores.
Despite growing interest in clinical LLMs, rigorous evaluation remains a fundamental bottleneck due to the scarcity of expert-annotated datasets. High-quality ground truth labels for complex clinical reasoning tasks require substantial investment in specialist time and expertise, limiting the scale at which models can be validated (Zac2023; Liang2017). Common evaluation approaches face significant limitations: synthetic data generation often fails to capture clinical complexity and is vulnerable to distributional shift (Santangelo2025; Yan2022; Chen2024; TIMER), while using actual patient treatments as ground truth is problematic, since real-world decisions frequently incorporate factors exogenous to guideline recommendations, such as patient preferences, drug availability, institutional protocols, and physician experience (Arts2016; Quaglini2004).
@@
Figure 1: CancerGUIDE framework. (1) Clinicians derive patient pathways ($label_{doc}$) from NCCN guidelines and real notes. These serve as gold-standard references to compare with LLM-derived pathways ($label_{LLM}$), producing a reference accuracy score ($Score_{r}$). (2) NCCN guidelines and both synthetic and real clinical notes are used to generate weak labels ($label_{proxy}$). LLM predictions ($label_{LLM}$) are compared to these proxy labels to compute proxy performance scores ($Score_{p}$), enabling evaluation of how well synthetic supervision approximates expert annotations. (3) Model consistency features (Feature Sets) and $Score_{p}$ are used to train a logistic regression meta-classifier that predicts whether a treatment recommendation is likely correct. This classifier is fit on labelled data from step 1 and unlabelled features from step 2. The classifier is applied at test time to accept or reject LLM-generated recommendations, supporting confidence estimation and threshold selection for clinical deployment.
@@
Our evaluation highlights the role of synthetic data and consistency-based data in proxy benchmarking, with these approaches achieving high Spearman correlation coefficients (r=0.88) and low RMSE values (RMSE=0.08). We demonstrate that model consistency can also serve as a reliable accuracy predictor, enabling our meta-learning framework to achieve an average of 0.804 AUROC in classifying individual prediction accuracy across all models, while unsupervised clustering using consistency signals alone achieves 0.666 F1 at separating correct from incorrect predictions.
@@
    Consistency framework for reliable treatment prediction: A hybrid agent and meta-classifier framework that generates relevant guideline paths and produces confidence values correlated with accuracy, enabling ROC curve calculation to clearly convey performance to clinicians and ensure compliance with regulatory standards (fda).
@@ 2.1 Guideline Adherent Treatment Recommendations
Guideline-adherent treatments have consistently been associated with improved clinical outcomes, such as overall survival (Dholakia2020; Lindqvist2022; Wckel2010). Clinical decision-support systems have the ability to translate clinical guidelines into point-of-care recommendations. Several studies have demonstrated the use of AI-driven systems in predicting guideline-adherent treatment recommendations in oncology (Nafees2023; Suwanvecho2021). Emerging work has also explored the utility of LLMs; for example, preliminary investigations suggest that ChatGPT can effectively summarize guideline content (Hamed2023) and exhibits partial concordance in identifying guideline-adherent treatments (Schulte2023). Furthermore, recent benchmarking efforts have systematically evaluated the alignment of LLMs with established medical guidelines, underscoring both their potential utility and the current limitations imposed by the relatively small number of verified clinical cases available (Fast2024).
@@ 2.2 Synthetic Generation of Clinical Data
Synthetic data generation is a promising strategy to address the challenges of privacy and data scarcity in healthcare. Past works have focused on generating high-fidelity electronic health records (EHRs) using generative models, such as generative adversarial networks, variational autoencoders and autoregressive models (Choi2017; Lu2022; Li2023; Yoon2023; Jarret2021). More recent frameworks have used LLMs to generate synthetic data and have shown improvements in privacy guarantees and scalability (Altalla2025; Barr2025; Taloni2025; Litake2024). In parallel, standardized evaluation metrics and scorecards have been proposed to assess the fidelity, privacy, and clinical utility of synthetic health data, providing a foundation for more rigorous benchmarking and deployment (Santangelo2025; Yan2022; Chen2024; medhelm). In addition, synthetic cohorts have been developed to model guideline-adherent treatment pathways, such as synthetic stroke registries for adherence benchmarking (Wang2025) and synthetic EHR modules that embed clinical practice guidelines and protocols (Walonoski2017). Previous work has explored guideline-following capability across frontier models on synthetic clinical data (li2025medguide). These efforts demonstrate the potential of synthetic data for adherence-focused tasks. However, significant challenges remain in generating high-fidelity, guideline-adherent cases while maintaining performance on downstream evaluation tasks.
@@ 2.3 Weak Supervision and Consistency
Weak supervision and consistency-driven learning have emerged as key strategies for training models when labeled data is limited or noisy. Weak supervision techniques, such as programmatic labeling and distant supervision, integrate heterogeneous, partially labeled sources to produce probabilistic labels for downstream predictive tasks (Mintz2009; Ratner2016; Ratner2017; Hsu2025; Falcon2025). Consistency-based methods encourage models to produce stable predictions under input perturbations, data augmentations, or repeated evaluations, helping to regularize training and improve generalization (Laine2016; Miyato2019; Xie2020; monkeys). Cycle consistency is a related method for generating weakly supervised labels from one data source. It involves transforming data into labels and then mapping those labels back to the original data to measure how well the label preserves the original information, providing a confidence estimate for the weakly supervised label (cycleconsistency). Pseudo-labeling, a related strategy, generates labels for unlabeled or partially labeled data by treating confident model predictions as additional supervised training data, enhancing model performance by expanding the training set with its own high-confidence predictions (Lee2013; Zou2021; Xie2020). Within healthcare, pseudo-labeling has shown to improve the reliability of EHR phenotyping and imaging tasks under label scarcity (Dong2023; Zhang2022; Nogues2022). Together, these techniques provide a framework for leveraging limited labeled data while maintaining prediction stability and adherence to domain-specific constraints, which is particularly relevant for guideline-adherent treatment modeling.
@@ 2.4 Non-Verifiable Task Evaluation
Evaluating LLMs on open-ended or non-verifiable tasks, such as treatment recommendation prediction, is challenging due to the scarcity of ground truth labels. Traditional metrics like accuracy or BLEU are insufficient in zero-label settings, prompting the use of human-aligned and/or proxy evaluation frameworks (Lyu2024; Yamauchi2025; Tan2024). Further, studies have shown that consistency-derived benchmarks provide improved predictive fidelity relative to ground-truth outcomes (Lee2024; Li2025). Proxy frameworks and consistency-derived benchmarks offer partial solutions, but their alignment with expert judgment remains imperfect, motivating the continued development of hybrid evaluation strategies.
@@ 3.1 Guideline-Compliance Task Formalization
To ground this formalization in clinical practice, we leveraged o3’s vision capabilities (openai2025o3) to extract the NCCN NSCLC guideline decision tree (NCCN) and curated an expert-annotated dataset. Oncologists traced patient notes through the decision tree, recording node sequences and assessing guideline adherence for each case. Further details on annotation procedures, participant recruitment, and quality control measures are provided in Appendix A.
@@
where $\mathcal{D}$ denotes the distribution over patients and $\mathcal{L}$ is standard loss. This formulation enables evaluation of $f$ in settings with limited or no access to human-labeled $(x,y)$ pairs, by leveraging model agreement, self-consistency, and benchmark proxy-derived features as signals for meta-classification.
@@ 3.1.1 Evaluation Preliminaries
Path Overlap captures the model’s ability to navigate the guideline-adherent decision process even if the final treatment is not fully correct, reflecting potential utility and acceptability in assisting clinicians. Treatment Match provides a direct measure of clinical accuracy for the final recommendation. Together, these metrics offer complementary perspectives: one emphasizes process fidelity and decision support potential, while the other captures correctness of the outcome. To ensure consistent and deterministic scores for path quality, LLM judges are not used to assess similarity of paths as this introduces stochasticity between models that confounds direct comparison in this setting.
@@ 3.2.1 Synthetic Supervision
Two complementary strategies are used: Structured Generation fills empty structured fields conditioned on the generated path and full clinical guidelines, performs consistency checks by reconstructing implied paths (discarding mismatched cases), then generates unstructured notes from structured data, target path, guidelines, and real clinical note examples. Unstructured Generation bypasses structured fields, generating synthetic notes directly from target paths, guidelines, and clinical note examples. Synthetic datasets were generated with GPT-4.1, except for those used to evaluate GPT-4.1, which were produced by GPT-5 under minimal-reasoning conditions.
@@ 3.2.2 Consistency-Based Pseudo-Labeling
[…]
@@
We further assign pseudo-labels by comparing predictions across the whole set of models, $M$. For a note $x$, if two or more models converge on the same pseudo-label after $k$ independent samples, we define this as the aggregated label. Convergence in this setting is exact path match. Inconsistent cases are simply excluded rather than penalized in contrast to the Self-Consistency method. We obtain the subset of (x, $\hat{y}$) pairs with the aggregated labels and evaluate all models in $M$ on this proxy benchmark.
@@ 3.3 Final Treatment Accuracy Prediction
We use consistency signals and model performance on proxy benchmarks to predict accuracy of a generated final treatment prediction. We train a meta-classifier using features derived from self-consistency metrics ($k$-rollout path overlap and treatment match), cross-model consistency (fraction of models with the same generated path), and proxy benchmark scores (synthetic and consistency-derived).
@@
To demonstrate label-free evaluation capabilities, we apply unsupervised methods to both accuracy classification and error identification. Clustering self- and cross-model consistency features naturally separates high-confidence correct from low-confidence incorrect predictions. For error identification, we tabulate inconsistencies across $k$ rollouts for each model $m$ on patient $p$. This approach leverages consistency signals to detect potential errors without relying on human labels.